OpenStudy (anonymous):

The sequence is geometric. What is the next term? 12, 24, 48, 96,...

OpenStudy (anonymous):

your common ration is 2, so you would multiply 96 by 2 to get 192

OpenStudy (anonymous):

okay so its 192 ?

OpenStudy (anonymous):

then what

jimthompson5910 (jim_thompson5910):

to find the common ratio, you pick any term (say 48) and divide it by the previous term (24) to get 48/24 = 2

OpenStudy (anonymous):

so my guess is an = -1(3)n-1 but im not sure

OpenStudy (anonymous):

i might have done it wrong

OpenStudy (anonymous):

wait my bad. You would use geometric sequence formula to plug in

OpenStudy (anonymous):

am i close?

OpenStudy (anonymous):

because my answer was an = -1(3)n-1

OpenStudy (anonymous):

uhh i think I'm using a different formula? I used \[a_n=a_1r^n\]

OpenStudy (anonymous):

so a1=12 because that's your first term and r is your ratio.

OpenStudy (anonymous):

yeah i just didnt know how to put the n below the a and the n-1 above

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

an = 3(-1)^n-1 ???

OpenStudy (anonymous):

????

jimthompson5910 (jim_thompson5910):

allie22 what's the sequence again?

OpenStudy (anonymous):

12, 24, 48, 96,...

jimthompson5910 (jim_thompson5910):

what is the first term

OpenStudy (anonymous):

12

jimthompson5910 (jim_thompson5910):

what is the common ratio

OpenStudy (anonymous):

;O idk!

jimthompson5910 (jim_thompson5910):

pick any term and divide it by the previous term

jimthompson5910 (jim_thompson5910):

that will get you the common ratio

OpenStudy (anonymous):

6

OpenStudy (anonymous):

idk i divided it by 2, this is confusing lol

jimthompson5910 (jim_thompson5910):

say you picked the second term 24 divide it by the previous term 12 to get 24/12 = 2

jimthompson5910 (jim_thompson5910):

the common ratio is 2

OpenStudy (anonymous):

ohhh okay

jimthompson5910 (jim_thompson5910):

first term: a = 12 common ratio: r = 2

jimthompson5910 (jim_thompson5910):

so that means \[\large a_{n} = a*(r)^{n-1}\] \[\large a_{n} = 12*(2)^{n-1}\] is the nth term formula

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

plug in n = 1 and tell me what you get

OpenStudy (anonymous):

an = 3(-1)^n-1

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

how are you getting that?

jimthompson5910 (jim_thompson5910):

that's something completely different

OpenStudy (anonymous):

wait let me try again

OpenStudy (anonymous):

186 ?

jimthompson5910 (jim_thompson5910):

what exactly are you looking for again?

OpenStudy (anonymous):

no wait 288

OpenStudy (anonymous):

what the next term is

OpenStudy (anonymous):

in the sequence

jimthompson5910 (jim_thompson5910):

oh all you do is multiply the last term by the common ratio 2

OpenStudy (anonymous):

192

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

but Rainie answered that already, so not sure where you're stuck